OpenStudy (anonymous):

please help (4,-6); m= 3/5

OpenStudy (amilapsn):

Do you want the equation of the straight line?

OpenStudy (anonymous):

thats the possable out comes.

OpenStudy (amilapsn):

\(m\) is the gradient of the line, right? How do you find the gradient of a straight line?

OpenStudy (anonymous):

idk

OpenStudy (triciaal):

use the same formula for the other question the gradient is the same as the slope

OpenStudy (anonymous):

its D?

OpenStudy (amilapsn):

\(\Huge m={\frac{\text{y difference}}{\text{x difference}}}\)

OpenStudy (anonymous):

ya

OpenStudy (amilapsn):

nope...

OpenStudy (amilapsn):

so, \(\Huge m=\frac{y-(-6)}{x-4}\) ok?

OpenStudy (anonymous):

ok

OpenStudy (amilapsn):

So what's the answer?

OpenStudy (anonymous):

-4?

OpenStudy (amilapsn):

I mean what's the option?

OpenStudy (anonymous):

Oh A

OpenStudy (amilapsn):

\[\Huge \cal Bingo!\]

OpenStudy (anonymous):

lol can you help on one more?

OpenStudy (amilapsn):

Yes sure.

OpenStudy (amilapsn):

What is the slope of the graph?

OpenStudy (anonymous):

how do i find that?

OpenStudy (amilapsn):

take two points of the graph and plugin the values to \[\Huge m={\frac{\text{y difference}}{\text{x difference}}}\]

OpenStudy (anonymous):

the 2 points are 10 and - 10 right??

OpenStudy (amilapsn):

Not quite. If you choose a point then that point has two coordinates: (x,y) right?

OpenStudy (anonymous):

ya

OpenStudy (amilapsn):

So?

OpenStudy (anonymous):

thats 10 and -10 thats the x,y

OpenStudy (anonymous):

10 and -2?

OpenStudy (amilapsn):

It's one point. To calculate slope you've to choose two points: Let's take point A and B. Can you tell me what's the coordinates of point A?

OpenStudy (anonymous):

4,-8?

OpenStudy (amilapsn):

nope. The coordinates of point A is (-8,6).. make sure to use brackets when writing coordinates. :) Can you try the coordinates of point B?

OpenStudy (anonymous):

-6,4

OpenStudy (amilapsn):

when you telling coordinates first you have to tell x coordinate. then you've to tell y coordinates.

OpenStudy (anonymous):

so 4,-6?

OpenStudy (amilapsn):

you got that. So what would be the slope of the graph?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i think i got it is it C?

OpenStudy (amilapsn):

Let's put the equation of the slope this way: \[\huge m=\frac {y_A-y_B}{x_A-x_B}\] Can you plugin the values?

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

HELLO?!

OpenStudy (amilapsn):

yep.... Do you need more explanation?

OpenStudy (anonymous):

no